OpenStudy (anonymous):

What is the solution to the following system using Cramer’s Rule? 2x + 7y = 2 3x + 8y = –2

OpenStudy (anonymous):

nevermind were good

OpenStudy (anonymous):

okay a1=2 and b1=7 c1=2....................a2=3 and b2=8 and c2=-2..........\[ D=\left[\begin{matrix}a1 & b1 \\ a2 & b2\end{matrix}\right] =a1b2-a2b1=2*8-3*7=-5\]

OpenStudy (anonymous):

I'm so lost..

OpenStudy (anonymous):

\[Dx=\left[\begin{matrix}c1 & b1 \\ c2 & b2\end{matrix}\right] =c1b2-c2b1=2*8--2*7=30\]

OpenStudy (anonymous):

the letters a1 b1 c1 are the coefficients the numbers that are in front of the x y and z, cramers rule involves matrices and determinants

OpenStudy (anonymous):

\[Dy=\left[\begin{matrix}a1 & c1 \\ a2 & c2\end{matrix}\right] =a1c2-a2c1=2*-2-3*2=-10\]

OpenStudy (anonymous):

the solution x=Dx/D =30/-5=-6 and y=Dy/D=-10/-5=2

OpenStudy (amistre64):

2x + 7y = 2 3x + 8y = –2 y = -2(2)-3(2) / 8(2)-3(7) x = -2(7)-8(2) / 3(7) - 8(2)

OpenStudy (amistre64):

well that inserted in the wrong place altogether :)